OpenStudy (anonymous):

cn anyone help me by answerg logarthmic diffrenttion : y =(x+1)^10 / (2x-4)^8 in term of log.. anyone, plss~

sam (.sam.):

Hmm lets see, you could try this, ln both sides \[\ln(y)=\ln[\frac{(x+1)^{10}}{(2x-4)^8}]\] \[\ln(y)=\ln[(x+1)^{10}]-\ln[(2x-4)^8]\] \[\ln(y)=10\ln[(x+1)]-8\ln[(2x-4)]\] Differentiate the terms individually